OpenStudy (anonymous):

Use the distributive property to simplify the expression below -5x+20÷-5

OpenStudy (anonymous):

−x − 4 x + 20 x + 4 x − 4

OpenStudy (anonymous):

possible answers

OpenStudy (campbell_st):

ok. so the question is \[\frac{-5x + 20}{-5}\] which can be written as \[\frac{-5x}{-5} + \frac{20}{-5} = \] just simplify the fractions...

OpenStudy (anonymous):

how?

OpenStudy (ja1):

@destineec Please refrain from giving out answers, if you can feel free to GUIDE the user to the right answer, not just blurting it out especially when someone is trying to help :) http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

was that even right?

OpenStudy (anonymous):

Oh okaay

OpenStudy (campbell_st):

you need to be able to cancel common factors... so look at it this way \[\frac{-5x + 20}{-5} = \frac{-5 \times x}{-5} + \frac{-5 \times -4}{-5}\] when the numerators are written in expanded form, what do you see as a common factor? cancel it and what are you left with

OpenStudy (skullpatrol):

$$\frac{bc}{bd}=\frac{c}{d}$$$$b\ne0,d\ne0$$